# Measuring cosmic baryon density with FRB and GW data

### v1

@@ Abstract
The dispersion measure (DM) analysis of extragalactic fast radio bursts (FRBs) has accounted for all cosmic baryons but remains limited by systematic uncertainties from various parameter degeneracies. We show that the prominent degeneracy between the baryon density ($\Omega_{\rm b}$) and the Hubble constant ($H_{0}$) can be disentangled through an independent $H_{0}$ value uniquely inferred from the absolute luminosity distance measurements of gravitational-wave (GW) standard sirens. By combining $104$ localized FRBs with $47$ GW events, we obtain a robust, late-Universe measurement of $\Omega_{\rm b}~{}$$=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN. Notably, incorporating GW data helps not only avoid biases induced by the $H_{0}$ tension, but also mitigate those from the parameters of diffuse baryon fraction and DM distribution models. Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
Cosmological parameters (339); Baryon density (139); Radio bursts (1339); Gravitational waves (678)
1 Introduction
@@
The concordance cosmology predicts consistency among independent probes across all epochs (Ostriker & Steinhardt, 1995). However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe (Cooke et al., 2018), about $30\%$ baryons remain undetected in the local Universe (Shull et al., 2012). These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” (Fukugita et al., 1998; Cen & Ostriker, 1999).
@@
Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin (Lorimer et al., 2007; Thornton et al., 2013). Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) (McQuinn, 2013). Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons (Deng & Zhang, 2014; Connor et al., 2025). Macquart et al. (2020) (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem. Yang et al. (2022) (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census. With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe (Zhou et al., 2014; Gao et al., 2014); see Bhandari & Flynn (2021),Wu & Wang (2024), and Glowacki & Lee (2025) for recent reviews.
@@
However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties (Kumar & Linder, 2019). The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq. (1)). This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM. Macquart20 modeled their probability distributions and Zhang et al. (2020, 2021) further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations. However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control. This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, the Hubble constant ($H_{0}$), and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq. (3)). Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature (Li et al., 2020; Qiang & Wei, 2020; Wang & Wei, 2023; Liu et al., 2025). Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB (Aghanim et al., 2020) (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) (Riess et al., 2022)) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the “$H_{0}$ tension” (Guo et al., 2019; Di Valentino et al., 2021; Abdalla et al., 2022; Hu & Wang, 2023). This inconsistency would introduce systematic uncertainties into the estimation of $\Omega_{\rm b}$ using FRB samples. It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 6767676774747474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Hagstotz et al. (2022)). Alternatively, jointly analyzing FRBs and datasets that independently measure $H_{0}$ would also be welcome.
@@
As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ (Schutz, 1986). This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis. Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known (Dalal et al., 2006; Cutler & Holz, 2009; Zhao et al., 2011; Cai & Yang, 2017; Abbott et al., 2017a, b, c; Chen et al., 2018; Cai et al., 2018; Wang et al., 2018; Zhang et al., 2019a, b; Zhang, 2019; Wang et al., 2020; Zhao et al., 2020b; Gray et al., 2020; Jin et al., 2020; Yu et al., 2020; Wang et al., 2022; Zhu et al., 2022a, b; Wu et al., 2023; Muttoni et al., 2023; Yun et al., 2023; Yu et al., 2024; Han et al., 2024; Feng et al., 2024; Song et al., 2024; Jin et al., 2024a, b; Li et al., 2024; Dong et al., 2024; Xiao et al., 2024; Zhu et al., 2024; Zheng et al., 2024; Han et al., 2025). During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and collected in the third Gravitational-Wave Transient Catalog (GWTC-3). Using 47 high–signal-to-noise ratio (SNR) GW standard sirens from GWTC-3, Abbott et al. (2023a) obtained a $\sim 10\%$ precision on $H_{0}$. Song et al. (2025) further combined these 47 GW standard sirens with an additional strong lensing system and achieved a model-independent constraint on $H_{0}$ with a precision of $\sim 8\%$. Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology.
Unlike previous studies that combined FRBs with traditional observations such as SNe (Jaroszynski, 2019; Gao et al., 2023) or CMB (Zhao et al., 2020a; Qiu et al., 2022), we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics. More recently, Zhang & Zhang (2025) and Gao et al. (2025b) proposed a novel method based on FRB/persistent radio sources (PRS) associations, which awaits the detection of more such events. The FRB + GW synergy is promising, with numerous detections expected from both communities in next-generation surveys (Zhang et al., 2023).
@@
In this Letter, we combine $104$ localized FRBs with $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$, without relying on any cosmological priors. To control the aforementioned systematics, we allow all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$). The parameters are defined in Sec. 2.1. to vary freely and incorporate GW data to break parameter degeneracies. Note that we do not assume any FRB/GW associations (see Wei et al. (2018) and Qiang et al. (2025) for studies based on such associations).
2 Data and methodology
2.1 Localized FRB data
@@
The average $\mathrm{DM}_{\mathrm{IGM}}$ in the standard $\Lambda$CDM (cosmological constant $\Lambda$ + cold dark matter) cosmology is given by the Macquart relation,
@@
where $\Delta\equiv{\rm{DM}}_{\rm{IGM}}/\left\langle{\rm{DM}}_{{{\rm{IGM}}}}\right\rangle$. $\alpha$ and $\beta$, which describe the inner gas density profile of halos, are best fitted as $\alpha=3$ and $\beta=3$ according to hydrodynamic simulations. The effective standard deviation $\sigma_{\mathrm{IGM}}=Fz^{-1/2}$ is parameterized by Macquart20, where $F$ is the parameter reflecting the galactic feedback. Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity. Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by (Zhang et al., 2025b) when normalizing Eq. (4).
@@
×$\times\exp\left[-\frac{(\ln\mathrm{DM}_{\rm{host}}-\mu_{\rm{host}})^{2}}{2\sigma_{\rm{host}}^{2}}\right],$ (5)
@@
For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (2.1):
@@
×$\times\int_{0}^{\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}}P_{\rm host}\left(\mathrm{DM}_{\rm host}\right)$
×$\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
@@
We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free. Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ (Zhang et al., 2021) and $P_{\rm host}(\mathrm{DM}_{\rm host})$ (Zhang et al., 2020). The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq. (4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq. (2.1), are interpolated from best-fit simulation results across redshift. Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table 2).
@@
Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ (Connor et al., 2025); (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\,\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments. The final sample is listed in Appendix B (see Table 2).
2.2 GW standard siren data

### v2

# Cosmic baryon census with fast radio bursts and gravitational waves

@@ Abstract
The cosmic baryon density fraction ($\Omega_{\rm b}$) is intrinsically correlated with the Hubble constant ($H_{0}$) through the critical density of the Universe. In the context of the decade-long $H_{0}$ tension, the significant discrepancy between early- and late-Universe measurements of $H_{0}$ implies that fixing its value or imposing an external prior could bias the baryon census. To address this concern, we construct a late-Universe probe framework that unifies fast radio bursts (FRBs) and gravitational-wave (GW) standard sirens, which can respectively resolve the “missing baryon” problem and the $H_{0}$ tension through their dispersion measures and absolute luminosity distances. By combining $104$ localized FRBs with $47$ GW events, we obtain an $H_{0}$-free measurement of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), in concordance with early-Universe observations of CMB + BBN. Although the current precision ($\sim 13\%$) is limited by sample size, the growing detections of both FRBs and GWs will make their synergy a powerful probe of low-redshift cosmology.
I Introduction
@@
The concordance cosmology predicts consistency among independent probes across all epochs Ostriker:1995su. However, while the cosmic microwave background (CMB) and Big Bang nucleosynthesis (BBN) provide precise estimates of the baryon content in the early Universe Cooke:2017cwo, about $30\%$ baryons remain undetected in the local Universe Shull:2011aa. These missing baryons are thought to reside in diffuse ionized gas that is too faint to be directly observed, known as the “missing baryon problem” Fukugita:1997bi; cen1999baryons.
@@
Fast radio bursts (FRBs) are bright, millisecond-duration radio transients of extragalactic origin lorimer2007; thornton2013pop, which can be broadly classified into repeaters and non-repeaters 2016Natur.531..202S; Sun2025a; Sun2025b. Their observed dispersion measures (DMs) trace the integrated column density of free electrons along the line of sight (LoS), making them sensitive to baryons in the intergalactic medium (IGM) mcquinn2013locating. Thus, FRBs with known redshifts derived from precise host galaxy localization can be used to trace the missing baryons Deng:2013aga; connor2025gas. Ref. Macquart:2020lln (hereafter Macquart20) first employed $5$ localized FRBs to estimate the present-day cosmic baryon density of $\Omega_{\rm b}=0.051^{+0.021}_{-0.020}$ at the $95\%$ confidence level, which agrees with the joint CMB and BBN result and likely resolves the local baryon problem. Ref. Yang:2022ftm (hereafter Yang22) used an updated sample of $22$ localized FRBs and obtained $\Omega_{\rm b}=0.0490_{-0.0033}^{+0.0036}$ at the $1\sigma$ level, reinforcing the evidence for a complete cosmic baryon census. With the high event rate and increasing number of localized events, FRBs have emerged as a promising cosmological probe Zhou:2014yta; Gao:2014iva; dai2023ska; Zhang:2023gye; see Refs. Bhandari:2021thi, Wu:2024iyu, and Glowacki:2024cgu for recent reviews.
@@
However, FRB cosmology based on DM is hindered by parameter degeneracies, which result in two main sources of systematic uncertainties Kumar:2019qhc. The first is from extracting the IGM contribution ($\mathrm{DM}_{\rm IGM}$) from the total DM budget (see Eq. (1)). This is mainly due to the difficulty in modeling both the host contribution ($\mathrm{DM}_{\rm host}$) and the LoS inhomogeneity in the IGM. Macquart20 modeled their probability distributions and Refs. Zhang:2020mgq; Zhang:2020xoc further employed the IllustrisTNG simulations to derive priors on these empirical DM model parameters for different FRB populations, which effectively reduce biases induced by host galaxy diversity and IGM fluctuations. However, the second uncertainty from the analysis of the extracted $\mathrm{DM}_{\rm IGM}$ remains challenging to control. This is because its redshift relation (known as the Macquart relation) exhibits intrinsic degeneracies among $\Omega_{\rm b}$, the Hubble constant ($H_{0}$), and the baryon fraction in diffuse ionized gas ($f_{\rm d}$) (see Eq. (3)). Given the limited constraints on $f_{\rm d}$, it is commonly modeled as either a constant or a redshift-dependent function in the literature Li:2020qei; Qiang:2020vta; Wang:2022ami; Liu:2025fdf. $\Omega_{\rm b}$ is intrinsically linked to $H_{0}$ through the critical density of the Universe. Although $H_{0}$ has been precisely measured, estimates from early- (e.g., Planck 2018 CMB 2020Planck (hereafter Planck18)) and late-Universe (e.g., SH0ES Cepheid-calibrated Type Ia supernovae (SNe) Riess:2021jrx) observations show a significant discrepancy now exceeding $5\sigma$, which has widely been known as the $H_{0}$ tension” Guo:2018ans; DiValentino:2021izs; Abdalla:2022yfr; Hu:2023jqc; CosmoVerseNetwork:2025alb. This inconsistency would introduce systematic uncertainties into the baryon census using FRB samples. It should be noted that Macquart20 and Yang22 derived their respective estimates of $\Omega_{\rm b}$ by assuming $H_{0}=70$ km s-1  Mpc-1and a flat prior over 67677474 km s-1  Mpc-1, which provides a reasonable strategy for handling the second uncertainty given the $H_{0}$ tension (note that many studies also constrain $H_{0}$ by adopting early-Universe priors on $\Omega_{\rm b}h^{2}$; see, e.g., Ref. Hagstotz:2021jzu). Nevertheless, any external assumption on $H_{0}$ could bias the baryon estimate, motivating the need for a direct inference of $\Omega_{\rm b}$ that is entirely untethered to the Hubble tension, e.g., jointly analyzing FRBs and datasets that independently measure $H_{0}$ in the late Universe.
@@
As an emerging cosmological probe, gravitational-wave (GW) standard sirens offer unique advantages for measuring $H_{0}$ Schutz:1986gp. This primarily arises from the fact that the amplitudes of GW waveforms are inversely proportional to the luminosity distances ($D_{\rm L}$) to the sources, allowing for a direct measurement of $D_{\rm L}$ through waveform analysis. Since $D_{\rm L}\propto 1/H_{0}$, GW standard sirens serve as a powerful and self-calibrating probe for determining $H_{0}$, provided that the redshifts of the GW sources are also known Dalal:2006qt; Cutler:2009qv; Zhao:2010sz; Cai:2016sby; LIGOScientific:2017adf; LIGOScientific:2017vwq; LIGOScientific:2017zic; Chen:2017rfc; Cai:2017aea; Wang:2018lun; Zhang:2019loq; Zhang:2018byx; Zhang:2019ylr; Wang:2019tto; Zhao:2019gyk; Gray:2019ksv; Jin:2020hmc; Yu:2020vyy; Mastrogiovanni:2021wsd; KAGRA:2021duu; Wang:2021srv; Zhu:2021aat; Zhu:2021bpp; Wu:2022dgy; Muttoni:2023prw; Yun:2023ygz; Yu:2023ico; Han:2023exn; Feng:2024lzh; Song:2022siz; Jin:2023sfc; Jin:2023tou; Li:2023gtu; Dong:2024bvw; Xiao:2024nmi; Zhu:2024qpp; Zheng:2024mbo; Han:2025fii; Du:2025odq; Song:2025bio; Jin:2025dvf; Dong:2025ikq. During the first three observing runs of the LIGO, Virgo, and KAGRA (LVK) collaboration, approximately 100 GW events were detected and compiled into the GWTC-3 catalog. Using 47 GW standard sirens from GWTC-3, Ref. LIGOScientific:2021aug achieved a $\sim 10\%$ precision on $H_{0}$, and Ref. Song:2025ddm further combined these sirens with an additional strong-lensing system to obtain a model-independent $\sim 8\%$ constraint. More recently, with the release of the LVK O4a data set, over 300 GW events have been reported, and by analyzing 142 of them while allowing the population-model parameters of both binary black holes (BBHs) and binary neutron stars (BNSs) to vary freely, the LVK collaboration has obtained a $\sim 14\%$ constraint on $H_{0}$ LIGOScientific:2025jau. Therefore, the $H_{0}$ information uniquely provided by GW data enables a direct estimate of $\Omega_{\rm b}$ disentangled from $H_{0}$ in the Macquart relation, which offers new insight into resolving systematic effects in FRB cosmology. Unlike previous studies that combined FRBs with traditional observations such as SNe Jaroszynski:2018vgh; gjz2023 or CMB Zhao:2020ole; Qiu:2021cww, we highlight that combining with GWs inherently involves calibration-free distances and avoids reliance on early-time physics, thereby eliminating the bias induced by the $H_{0}$ tension.
@@
In this work, we present the first $H_{0}$-free baryon census at low redshift, within a late-Universe probe framework that unifies $104$ localized FRBs and $47$ GW events from the GWTC-3 catalog to jointly constrain $\Omega_{\rm b}$ and $H_{0}$. To control model-dependent systematics, we impose no FRB population model and fully marginalize over all DM parameters [^1] In this context, we refer to the DM parameters as ($f_{\rm d}$, $F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$), while the empirical DM model parameters refer specifically to ($F$, $\mu_{\rm host}$, and $\sigma_{\rm host}$). The parameters are defined in Sec. II.1., while on the GW side we adopt a population model for the main analysis and additionally release its parameters to assess robustness. Note that we do not assume any FRB/GW associations (see Refs. Wei:2018cgd and Qiang:2024erm for studies based on such associations).
This work is organized as follows. In Sec. II, we describe the datasets and likelihood construction for both FRBs and GWs. Results and related discussions are presented in Sec. III, and we conclude in Sec. IV.
II Data and methodology
II.1 Localized FRB data
@@
[…]
@@
where $\Delta\equiv{\rm{DM}}_{\rm{IGM}}/\left\langle{\rm{DM}}_{{{\rm{IGM}}}}\right\rangle$. $\alpha$ and $\beta$, which describe the inner gas density profile of halos, are best fitted as $\alpha=3$ and $\beta=3$ according to hydrodynamic simulations. The effective standard deviation $\sigma_{\mathrm{IGM}}=Fz^{-1/2}$ is parameterized by Macquart20, where $F$ is the parameter reflecting the galactic feedback. Here, $A$ is a normalization constant, and $C_{0}$ is chosen to ensure that the mean of this distribution is unity. Note that we include the factor $1/\langle{\rm DM_{\rm IGM}}\rangle$ suggested by Ref.  Zhang:2025wif when normalizing Eq. (4).
@@
$\times\exp\left[-\frac{(\ln\mathrm{DM}_{\rm{host}}-\mu_{\rm{host}})^{2}}{2\sigma_{\rm{host}}^{2}}\right],$ (5)
@@
For the $i$-th FRB, the possibility distribution of its extragalactic DM component ($\mathrm{DM}^{\prime}=\rm DM_{\rm IGM}+\rm DM_{\rm host}$) is computed by convolving Eqs. (2), (4), and (II.1):
@@
$\times\int_{0}^{\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}}P_{\rm host}\left(\mathrm{DM}_{\rm host}\right)$
$\times P_{\rm IGM}\left(\mathrm{DM}^{\prime}_{i}-\mathrm{DM}_{\rm halo}-\mathrm{DM}_{\rm host}\right)\,\mathrm{dDM}_{\rm host}.$ (7)
@@
We also consider a scenario in which the empirical DM model parameters are fixed, though our main results treat them as free. Specifically, we adopt values calibrated from cosmological hydrodynamical simulations, notably the IllustrisTNG simulation, which provides redshift-dependent distributions for both $P_{\rm IGM}(\mathrm{DM}_{\rm IGM})$ Zhang:2020xoc and $P_{\rm host}(\mathrm{DM}_{\rm host})$ Zhang:2020mgq. The parameters $A(z)$, $C_{0}(z)$, and $\sigma_{\rm IGM}(z)$ in Eq. (4), as well as $\mu_{\rm host}(z)$ and $\sigma_{\rm host}(z)$ in Eq. (II.1), are interpolated from best-fit simulation results across redshift. Note that $\mu_{\rm host}$ and $\sigma_{\rm host}$ are further fixed based on the host galaxy types (see the seventh column of Table LABEL:tab:localized_FRBs).
@@
Finally, we select a sample of $104$ localized FRBs from an initial set of $119$, based on the following criteria: (1) $\mathrm{DM}_{\mathrm{MW,ISM}}$ is required to be less than $40\%$ of $\mathrm{DM}_{\rm obs}$ connor2025gas; (2) $\mathrm{DM}_{\rm obs}-\rm DM_{MW,ISM}$ is taken to exceed 80 $\rm{pc}\,\rm{cm}^{-3}\,$, ensuring a positive $\mathrm{DM}_{\rm host}$; (3) FRB 20190520B and FRB 20220831A are excluded due to excess local DM contributions from their host environments. The final sample is listed in Appendix B (see Tables LABEL:tab:localized_FRBs).
II.2 GW standard siren data